# The Great Comets of 1843 and 1882 at Their Previous Return to Perihelion in the Twelfth Century: One Spectacular, the Other Dull

### v1

@@ Abstract
New insights into the history of C/1843 D1 and C/1882 R1, the two celebrated Kreutz sungrazers, are provided by assessing evidence on their appearance at the previous perihelion return, known as X/1106 C1 and the Chinese comet of 1138 (Ho’s No. 403), respectively. The conditions differed vastly because of disparities in geocentric distance, solar elongation, and phase correction (forward scattering), all linked to the arrival times (early February vs early August). The conclusions include: the daytime sighting of the 1106 comet by Sigebert de Gembloux is consistent with expectation and so are the accounts of an exceptionally long tail observed later in twilight; the comet reached perihelion only hours before its daytime detection; the 1138 comet could have never been sighted in daylight or discovered much earlier than it actually was, in early September, one month after perihelion; at discovery, the tail is predicted to have reached elevations of ${}^{\circ}\!$, while the head was merely $\sim$10 above horizon, when observed from moderate northern latitudes; notion that Kreutz sungrazers at perihelion between mid-May and mid-August could not be seen from the ground except possibly in daylight is misleading; the appearance of the 1138 comet’s nucleus after its tidal fragmentation at perihelion is modeled on the assumption that it consisted of five major fragments (including C/1882 R1 and C/1965 S1); and unpredictable morphological changes with time in the 1882 sungrazer’s split nucleus are discussed.
@@ 1. Introduction
The bottom line is that the 1843 and 1882 sungrazers must have very similar orbital periods, having passed their previous perihelion in the early 12th century in the same order and at times closer to one another than in the 19th century. Since only one suspect in the early 12th century had until recently been known — the Great Comet of 1106 (X/1106 C1) — we were confronted with the problem of a missing second sungrazer.
@@ 2. Search for the Missing Sungrazer:
One of the issues that I am having with this widely accepted notion involves the comet of 1041 in Hasegawa & Nakano’s (2001) list of suspected sungrazers. This object, observed in China and Korea from September 1 on, was estimated by the authors to have passed perihelion on August 4 $\pm$ 4, more than a week before the end of the disadvantaged time zone. No daylight sightings were reported, but the comet was observed about four weeks after perihelion, when Hasegawa & Nakano computed its location to have been some 1.7 AU from the Earth, 1 AU from the Sun, and a solar elongation of 35.
When Sekanina & Kracht’s (2022) in-depth investigation of the motion of Ikeya-Seki (C/1965 S1) showed that the comet previously was at perihelion around the year 1140 and the Great Comet of 1106 (X/1106 C1) could not be related to either Ikeya-Seki or the Great September Comet of 1882, perfunctory search in Ho’s (1962) catalogue resulted in the discovery of a Chinese comet of 1138, under No. 403, as the likely candidate. Entirely by coincidence, its timeline strongly resembled the timeline of the comet of 1041: the first observation on September 3 and an estimated perihelion time on August 1.
@@
Although the comet of 1041 is of major interest as a possible previous appearance of comet Pereyra (Sekanina & Kracht 2022), the primary subject of this paper is the Chinese comet of 1138 as the previous appearance of the Great September Comet of 1882 and comet Ikeya-Seki. In the following I compare the 1138 comet with the Great Comet of 1106; this is particularly appropriate because both objects are believed to have been of about equal intrinsic brightness, yet the 1106 comet was seen in broad daylight $\sim$1 from the Sun on February 2 and displayed a spectacular tail up to $\sim$100 in the following weeks (e.g., Kronk 1999), while the 1138 comet was a lackluster object, reported for the first time more than 30 days after the estimated perihelion time. Yet, the 1106 comet disappeared some 45–50 days after perihelion, while the 1138 comet about 60 days after perihelion. This effect may have been a by-product of the 1138 comet’s more extensive perihelion fragmentation, which is believed to lead to a slower rate of post-perihelion fading (Sekanina 2022; Section 3.1).
@@ 3.1. Adopted Light Curves of the 1106 and 1138 Comets
The probable perihelion-fragmentation histories of the 1106 and 1138 comets were addressed in one of my recent papers (Sekanina 2025). I speculated that besides the Great March Comet of 1843, the main returning mass of the Great Comet of 1106, comet C/1668 E1 was a major preceding fragment and another large piece could arrive at perihelion around the year 2050. Their possible early sibling — a comet of 1529, and another one, which might arrive in the mid-23rd century, are just extremely remote possibilities. Adopting $\nu_{\rm frg}=3$, I get $9.52.5\,n^{+}\!=$ 9.5 from Equation (3). With the perihelion distance from Table 1, the absolute post-perihelion magnitude of the 1106 comet is $H_{0}^{+}\!=+1.7$ from Equation (2).
@@
The assumptions of an identical preperihelion light curve with the 1106 comet and a stronger propensity for fragmentation at perihelion leads to the 1138 comet being intrinsically brighter after perihelion, by more than 1 magnitude at 1 AU from the Sun.
@@ 3.3. Visibility of the Comets in Broad Daylight
The apparent light curves of the two comets and the respective limiting magnitudes are plotted in Figure 1. The striking feature is the reversed order of the light curves. Whereas Equations (4) showed the 1138 comet to be more than 1 mag intrinsically brighter after perihelion, Figure 1 predicts that the 1106 comet should have appeared distinctly brighter to the terrestrial observer. The driver of this reversal was the differential phase effect early after perihelion, while later the disparity in the geocentric distance became more important. And, as follows from Table 2, for a given heliocentric distance (which governs the intrinsic-brightness variations) the solar elongation of the 1138 comet was always only about one half of the solar elongation of the 1106 comet, a major factor affecting the visibility of the two objects.
@@
Figure 2.— Projection onto the plane of the sky of the predicted orbit for X/1106 C1 close to perihelion (equinox J2000). The dotted part of the orbit indicates that the comet was then behind the Sun. The bullets show the locations of the comet at the times, reckoned in hours from perihelion. Parenthesized are the respective phase angles. Also shown are arcs of solar elongations of ${}^{\circ}\!\!$.6, ${}^{\circ}\!$, and ${}^{\circ}\!\!$.6. Receding from the Sun at a nearly constant projected rate of $\sim$4 per day starting shortly after perihelion, the comet was observed at Gembloux as a bright star over a period of six hours.
@@ 4. Daytime Observation of X/1106 C1
The extremely favorable daylight observing conditions predicted for the 1106 comet appear to be confirmed by the historic record, written by Sigebert de Gembloux in his Chronicon sive Chronographia in AD $1$. A monk in the Benedictine Abbey of Gembloux, Belgium, he noted in the chronicle that on 1106 February 2 a star appeared during the daytime, between the third and ninth hour, about a cubit from the Sun. [^2] To answer the question of what was the distance in degrees is not straightforward. One can find a lot of information on the cubit in the literature, including the differences between the Sumerian cubit, Egyptian royal cubit, Greek cubit, Roman cubit, etc. and anthropologists still write scientific papers on the cubit in the 21st century (e.g., Stone 2014). Yet, all this information is irrelevant to Sigebert’s observation, because the primary usage of the term cubit in the antiquity was as a unit of length, equal to the extent of the forearm from the tip of the middle finger to the elbow. Different cubits had different lengths, mostly between 17.5 and 21 inches, or 44 and 53 cm. — Obviously, the cubit as a measure of length and the cubit as a measure of angular separation are two very different quantities, and it is unfortunate that for some strange reason the same term is used for both of them. It is more difficult to find information about the cubit measuring angular separation. And this information is not only scarce but, worse, also inconsistent. When trying to google the term, one usually finds that 1 cubit is approximately 1, with no details or reference being given. There are, however, exceptions. To an inputed request “cubit angular measure” Google’s AI Overview returns: A “cubit” as an angular measure roughly corresponded to a finger’s width held at arm’s-length, which subtends an angle of about 2 degrees. Similarly, in Wikipedia’s article on “Cubit” one can read: There is some evidence that cubits were used to measure angular separation. The Babylonian Astronomical Diary for 568–567 BCE refers to Jupiter being one cubit behind the elbow of Sagittarius. One cubit measures about 2 degrees, with a reference to Steele (2008), noting that the relationship between the cubit as a unit of length and a unit of angular separation was not in the book elaborated on. Interestingly, I was unable to find the cubit, as an angular measure, in Wikipedia’s article “List of obsolete units of measurement,” or among the angular measure units in Wikipedia’s article “List of unusual units of measurement,” or in any of the associated articles “Ancient Arabic (Egyptian, Greek, Mesopotamian, Roman) units of measurement.” (I have not investigated sources concerning the works by Otto E. Neugebauer, as this would involve prohibitive time demands.) On the other hand, in their paper on ancient Babylonian time measurements Stephenson & Fatoohi (2008) wrote that it should be noted that observed angular distances between celestial bodies were almost invariably expressed by the Babylonians in terms of the kùš (cubit), which was equivalent to between about 2 and 2.5 degrees. The various sources clearly provide a rather wide range of numbers. Using the defining ratio of the finger’s width to arm’s-length, taking the standard values of 1.85 cm for the finger’s width and $\frac{14}{9}$ for the ratio of arm’s-length to forearm’s-length of 46 cm, I derive that 1 cubit = ${}^{\circ}\!$.5, yet another estimate.
@@
In any case, to inspect the timeline more closely, I plot in Figure 2 the orbit of the 1106 comet in projection onto the plane of the sky from 3.5 hours before perihelion to nearly 13 hours after perihelion. For approximately one hour, the comet was hidden behind the Sun, including its passage through perihelion. When observed at the Abbey by Sigebert himself and/or his colleagues, the comet was already past perihelion and moving to the southeast of the Sun at a nearly constant rate of $\sim$10 per hour in the general direction of the Earth.
It turns out that from Sigebert’s description one can approximately establish the UT times of observation and therefore the comet’s perihelion time. Given that the geographic longitude and latitude of the Abbey were, respectively, ${}^{\circ}\!$.70 E and ${}^{\circ}\!$.56 N, and the elevation was 150 meters above sea level, the Sun rose at 7:13 UT and set at 16:42 UT on 1106 February 2 (not corrected for refraction). Daylight thus extended over 9:29 hours and, rounded to the nearest quarter of an hour, the comet was seen between 10:15 and 16:15 UT. Since the latter time was less than 30 minutes before sunset, it is reasonable to conclude that the event’s reported termination was a corollary to the deteriorating atmospheric conditions at low elevations, not the comet’s fading. During the six-hour observing window, the comet’s solar elongation, $\epsilon$, is predicted from Figure 2 to have increased by slightly more than 1. More precisely, the relationship between $\epsilon$ and the time from perihelion, $\tau$ (in days), is given by an empirical formula:
@@
with a mean error of ${}^{\circ}\!$.0044 or ${}^{\>\!\prime\!}$.26 and applicable to an interval of +0.18 day $\leq\!\tau\!\leq$ +0.80 day. For example, if 1 cubit is taken to equal ${}^{\circ}\!$.1, the comet’s solar elongation would have been ${}^{\circ}\!$.6 some 4.4 hours after perihelion and ${}^{\circ}\!$.6 about 5.7 hours later. The elongation would equal 1 cubit 7:15 hours after perihelion, equivalent to the middle of Sigebert’s six-hour long window, 13:15 UT. One finds in this case that the comet passed perihelion at 6:00 UT, or on February 2.25 UT.
@@
Figure 3.— Light curve (thick line), limiting magnitude (thin line), and solar elongation (dotted line) predicted for comet X/1106 C1 vs time on 1106 February 2, when the comet was observed as a bright star over a period of six hours in broad daylight at close proximity of the Sun, according to a historical record by Sigebert, a monk in the Benedictine Abbey of Gembloux, Belgium. The comet’s brightness, greatly enhanced by the effect of forward scattering (up to nearly 4 mag at maximum) is predicted to be some 4–6 magnitudes above the visibility threshold for the naked eye, except at times of very low elevation above the horizon, just after sunrise and before sunset. The start and end of the six-hour long period of observation are known with only limited accuracy. The comet was assumed to pass perihelion at 6:00 UT on February 2.
Figure 3 shows that the comet’s brightness stayed above the predicted visibility threshold by 4–6 magniudes except when very low above the horizon. However, it should be noted that the comet profited greatly from effects of forward scattering, which at peak augmented its brightness by nearly 4 magnitudes, about 3 hours after perihelion, and contributed more than 2.5 magnitudes throughout the period of time between 7:30 and 16:30 UT. Interestingly, the end of Sigebert’s observation coincides with the time when the comet’s light curve meets the limiting magnitude. The inaccuracies in Sigebert’s report affect choice of the perihelion time, which influences the positions of the three curves in Figure 3. For an earlier time, the comet’s predicted solar elongation would be greater and the comet would be fainter but the detection limit would drop as well. In any case, the comet must have offered an incredible sight and Sigebert, in retrospect, left a valuable piece of scientific data.

### v2

@@ Abstract
New insights into the history of C/1843 D1 and C/1882 R1, the two celebrated Kreutz sungrazers, are provided by assessing evidence on their appearance at the previous perihelion return, known as X/1106 C1 and the Chinese comet of 1138 (Ho’s No. 403), respectively. The conditions differed vastly because of disparities in geocentric distance, solar elongation, and phase correction (forward scattering), all linked to the arrival times (early February vs early August). The conclusions include: the daytime sighting of the 1106 comet by Sigebert de Gembloux is consistent with expectation and so are the accounts of an exceptionally long tail observed later in twilight; the comet reached perihelion only hours before its daytime detection; the 1138 comet could have never been sighted in daylight or discovered much earlier than it actually was, in early September, one month after perihelion; at discovery, the tail is predicted to have reached elevations of ${}^{\circ}\!$, while the head was only 10 above horizon, when observed from moderate northern latitudes; the notion that Kreutz sungrazers at perihelion between mid-May and mid-August could not be seen from the ground except possibly in daylight is misleading; the appearance of the 1138 comet’s nucleus after its tidal fragmentation at perihelion is modeled on the assumption that it consisted of five major fragments (including C/1882 R1 and C/1965 S1); and unpredictable morphological changes with time in the 1882 sungrazer’s split nucleus are discussed.
@@ 1. Introduction
The similarity of the orbital periods suggested that the 1843 and 1882 sungrazers had passed their previous perihelion, apparently in the early 12th century, in the same order and at times even closer to one another than in the 19th century. Since only one suspect in the early 12th century had until recently been known — the Great Comet of 1106 (X/1106 C1) — we were confronted with the problem of a missing second sungrazer.
@@ 2. Search for the Missing Sungrazer:
One of the issues that I am having with this widely accepted notion involves the comet of 1041 in Hasegawa & Nakano’s (2001) list of suspected sungrazers. This object, observed in China and Korea from September 1 on, was estimated by the authors to have passed perihelion on August 4 $\pm$ 4, more than a week before the end of the disadvantaged time zone. No daylight sightings were reported, but the comet was observed about four weeks after perihelion, when Hasegawa & Nakano computed its position to have been some 1.7 AU from the Earth, 1 AU from the Sun, and a solar elongation of 35.
When Sekanina & Kracht’s (2022) in-depth investigation of the motion of Ikeya-Seki (C/1965 S1) showed that the comet previously was at perihelion around the year 1140 and the Great Comet of 1106 (X/1106 C1) could not be the parent of either Ikeya-Seki or the Great September Comet of 1882, perfunctory search in Ho’s (1962) catalogue resulted in the discovery of a Chinese comet of 1138, under No. 403, as the likely candidate. Entirely by coincidence, its timeline strongly resembled the timeline of the comet of 1041: the first observation on September 3 and an estimated perihelion time on August 1.
@@
Although the comet of 1041 is of major interest as a possible previous appearance of comet Pereyra (Sekanina & Kracht 2022), the primary subject of this paper is the Chinese comet of 1138 as the previous appearance of the Great September Comet of 1882 and comet Ikeya-Seki. In the following I compare the 1138 comet with the Great Comet of 1106; this is particularly appropriate because both objects are believed to have been of about equal intrinsic brightness, yet the 1106 comet was seen in broad daylight $\sim$1 from the Sun on February 2 and displayed a spectacular tail up to $\sim$100 long in the following weeks (e.g., Kronk 1999), while the 1138 comet was a lackluster object, reported for the first time more than 30 days after the estimated perihelion time. Yet, the 1106 comet disappeared some 45–50 days after perihelion, while the 1138 comet about 60 days after perihelion. This effect may have been a by-product of the 1138 comet’s more extensive perihelion fragmentation, which is believed to lead to a slower rate of post-perihelion fading (Sekanina 2022; Section 3.1).
@@ 3.1. Adopted Light Curves of the 1106 and 1138 Comets
The possible perihelion-fragmentation histories of the 1106 and 1138 comets were addressed in one of my recent papers (Sekanina 2025). I speculated that besides the Great March Comet of 1843, the main returning mass of the Great Comet of 1106, comet C/1668 E1 was a major preceding fragment and another large piece could arrive at perihelion around the year 2050. Their possible early sibling — a comet of 1529, and another one, which might arrive in the mid-23rd century, are just two rather remote possibilities. Adopting $\nu_{\rm frg}=3+$, I get $2.5\,n^{+}\!\simeq 9.3$ from Equation (3). With the perihelion distance from Table 1, the absolute post-perihelion magnitude of the 1106 comet is $H_{0}^{+}\!=+1.2$ from Equation (2).
@@
The assumptions of an identical preperihelion light curve with the 1106 comet and a stronger propensity for fragmentation at perihelion leads to the 1138 comet being intrinsically brighter after perihelion, by nearly 1 magnitude at 1 AU from the Sun.
@@ 3.3. Visibility of the Comets in Broad Daylight
The apparent light curves of the two comets and the respective limiting magnitudes are plotted in Figure 1. The striking feature is the reversed order of the light curves. Whereas Equations (4) showed the 1138 comet to be almost 1 mag intrinsically brighter at 1 AU after perihelion, Figure 1 predicts that the 1106 comet should have appeared distinctly brighter to the terrestrial observer. The driver of this reversal was the differential phase effect early after perihelion, while later the disparity in the geocentric distance became more important. And, as follows from Table 2, for a given heliocentric distance (which governs the intrinsic-brightness variations) the solar elongation of the 1138 comet was always only about one half of the solar elongation of the 1106 comet, a major factor affecting the visibility of the two objects.
@@
Figure 2.— Projection onto the plane of the sky of the predicted orbit for X/1106 C1 close to perihelion (equinox J2000). The dotted part of the orbit indicates that the comet was then behind the Sun. The bullets show the locations of the comet at the times, reckoned in hours from perihelion. Parenthesized are the respective phase angles. Also shown are arcs of solar elongations of ${}^{\circ}\!\!$.6, ${}^{\circ}\!$, and ${}^{\circ}\!\!$.6. Receding from the Sun at a nearly constant projected rate of $\sim$4 per day starting shortly after perihelion, the comet was observed at Gembloux in daylight as a star over a period of six hours.
@@ 4. Daytime Observation of X/1106 C1
The extremely favorable daylight observing conditions predicted for the 1106 comet appear to have facilitated an amazing sighting, described by Sigebert de Gembloux in his Chronicon sive Chronographia in AD $1$. A monk in the Benedictine Abbey of Gembloux, Belgium, he noted in the chronicle that on 1106 February 2 a star appeared during the daytime, between the third and ninth hour, about a cubit from the Sun. [^2] To answer the question of what was the distance in degrees is not straightforward. One can find a lot of information on the cubit in the literature, including the differences between the Sumerian cubit, Egyptian royal cubit, Greek cubit, Roman cubit, etc. and anthropologists still write scientific papers on the cubit in the 21st century (e.g., Stone 2014). Yet, all this information is irrelevant to Sigebert’s observation, because the primary usage of the term cubit in the antiquity was as a unit of length, equal to the extent of the forearm from the tip of the middle finger to the elbow. Different cubits had different lengths, mostly between 17.5 and 21 inches, or 44 and 53 cm. — Obviously, the cubit as a measure of length and the cubit as a measure of angular separation are two very different quantities, and it is unfortunate that for some strange reason the same term is used for both of them. It is more difficult to find information about the cubit measuring angular separation. And this information is not only scarce but, worse, also inconsistent. When trying to google the term, one usually finds that 1 cubit is approximately 1, with no details or reference being given. There are, however, exceptions. To an inputed request “cubit angular measure” Google’s AI Overview returns: A “cubit” as an angular measure roughly corresponded to a finger’s width held at arm’s-length, which subtends an angle of about 2 degrees. Similarly, in Wikipedia’s article on “Cubit” one can read: There is some evidence that cubits were used to measure angular separation. The Babylonian Astronomical Diary for 568–567 BCE refers to Jupiter being one cubit behind the elbow of Sagittarius. One cubit measures about 2 degrees, with a reference to Steele (2008), noting that the relationship between the cubit as a unit of length and a unit of angular separation was not in the book elaborated on. Interestingly, I was unable to find the cubit, as an angular measure, in Wikipedia’s article “List of obsolete units of measurement,” or among the angular measure units in Wikipedia’s article “List of unusual units of measurement,” or in any of the associated articles “Ancient Arabic (Egyptian, Greek, Mesopotamian, Roman) units of measurement.” (I have not investigated sources concerning the works by Otto E. Neugebauer, as this would involve prohibitive time demands.) On the other hand, in their paper on ancient Babylonian time measurements Stephenson & Fatoohi (2008) wrote that it should be noted that observed angular distances between celestial bodies were almost invariably expressed by the Babylonians in terms of the kùš (cubit), which was equivalent to between about 2 and 2.5 degrees. The various sources clearly provide a rather wide range of numbers. Using the defining ratio of the finger’s width to arm’s-length, taking the standard values of 1.85 cm for the finger’s width and $\frac{14}{9}$ for the ratio of arm’s-length to forearm’s-length of 46 cm, I derive that 1 cubit = ${}^{\circ}\!$.5, yet another estimate.
@@
In any case, to inspect the timeline more closely, I plot in Figure 2 the orbit of the 1106 comet in projection onto the plane of the sky from 3.5 hours before perihelion to nearly 13 hours after perihelion. For approximately one hour, the comet was hidden behind the Sun, including its perihelion point. When observed at the Abbey by Sigebert himself and/or his colleagues, the comet was already past perihelion and moving to the southeast of the Sun at a nearly constant rate of $\sim$10 per hour in the general direction of the Earth.
It turns out that from Sigebert’s description one can approximately establish the UT times of observation and therefore the comet’s perihelion time. Given that the geographic longitude and latitude of the Abbey were, respectively, ${}^{\circ}\!$.70 E and ${}^{\circ}\!$.56 N, and the elevation was 150 meters above sea level, the Sun rose at 7:13 UT and set at 16:42 UT on 1106 February 2 (not corrected for refraction). Daylight thus extended over 9:29 hours and, rounded to the nearest quarter of an hour, the comet was seen between 10:15 and 16:15 UT. Since the latter time was less than 30 minutes before sunset, it is reasonable to conclude that the event’s reported termination was a corollary to the deteriorating atmospheric conditions at low elevations, not the comet’s sudden fading. During the six-hour observing window, the comet’s solar elongation, $\epsilon$, is predicted from Figure 2 to have increased by slightly more than 1. The relationship between $\epsilon$ and the time from perihelion, $\tau$ (in days), is given more accurately by an empirical formula:
@@
with a mean error of ${}^{\circ}\!$.0044 or ${}^{\>\!\prime\!}$.26 and applicable to an interval of +0.18 day $\leq\!\tau\!\leq$ +0.80 day. With the solar elongation of ${}^{\circ}\!$.6 about 4.4 hours after perihelion and ${}^{\circ}\!$.6 about 5.7 hours later, the comet would have reached an elongation of 1 cubit equaling ${}^{\circ}\!$.1 some 7:15 hours after perihelion, equivalent to 13:15 UT, the middle of Sigebert’s six-hour long window. One finds in this case that the comet had passed perihelion at 6:00 UT, or on February 2.25 UT.
@@
Figure 3.— Light curve (thick line), limiting magnitude (thin line), and solar elongation (dotted line) predicted for comet X/1106 C1 vs time on 1106 February 2, when the comet was observed as a star over a period of six hours in broad daylight at close proximity of the Sun, according to a historical record by Sigebert, a monk in the Benedictine Abbey of Gembloux, Belgium. The comet’s brightness, greatly enhanced by the effect of forward scattering (up to nearly 4 mag at maximum) is predicted to be some 4–6 magnitudes above the visibility threshold for the naked eye, except at times of very low elevation above the horizon, just after sunrise and before sunset. The start and end of the six-hour long period of observation are known with only limited accuracy. The comet was assumed to pass perihelion at 6:00 UT on February 2.
Figure 3 shows that the comet’s brightness stayed above the predicted visibility threshold by 4–6 magniudes except when very low above the horizon. However, it should be noted that the comet profited greatly from effects of forward scattering, which at peak augmented its brightness by nearly 4 magnitudes, about 3 hours after perihelion, and contributed more than 2.5 magnitudes throughout the period of time between 7:30 and 16:30 UT. Interestingly, the end of Sigebert’s observation coincides with the time when the comet’s light curve meets the limiting magnitude. Inaccuracies in Sigebert’s report affect choice of the perihelion time, which influences the positions of the three curves in Figure 3. For an earlier time, the comet’s predicted solar elongation would be greater and the comet would be fainter, but the visibility limit would drop as well. In any case, the comet must have offered an incredible sight and Sigebert, in retrospect, left a valuable piece of scientific data.